A squirrel has x- and y-coordinates (1.1 m, -0.5 m) at time t1 = 0 sec and coordinates (5.3 m, 3.4 m) at time t2 = 3.0 sec. For this interval, find (a) the components of average velocity; (b) the magnitude and direction of average velocity.


1
Expert's answer
2021-07-19T09:50:02-0400

a)


"v_x=\frac{5.3-1.1}{3}=1.4\frac{m}{s}\\v_y=\frac{3.4+0.5}{3}=1.3\frac{m}{s}"

b)


"v=\sqrt{1.4^2+1.3^2}=1.9\frac{m}{s}\\\theta=\arctan{\frac{1.3}{1.4}}=43\degree"